# thanks: These authors contributed equally to this work

### v1

@@ III.2 Heralded Purification Protocol- HPP
In Fig. 2(b), we present the experimental configuration of the heralded purification protocol (BB84+HPP). In this scheme, we consider a SPS consisting for example a gCQD on a hybrid nanocone-antenna device (titled here SPS2). Such a device showed both $\eta_{C}\simeq 1$ together with high attainable values of $P_{2}$, resulting from a large Purcell factor induced by the nanocone which significantly enhances $QY_{XX}$ [34, 44]. Alice operates at a single excitation power $S_{3}>1$, deep in the saturation regime, to excite the BX state with very high probability, thus maximizing $P_{2}$, which is now only limited by $QY_{XX}$ and $QY_{X}$ [34]. A beam-splitter (BS) and a single photon detector (SPD) are added as a purification stage after the SPS in Alice’s system, and before the standard BB84 encoding unit. As $P_{N>2}$ are negligible (see Supp. Sec. III), a real photon detection event in Alice’s detector sets $\tilde{P}_{2}=0$ for that pulse, where $\tilde{P}_{n}$ is the effective photon number probability after the purification stage. In the sifting step of the QKD protocol, only events with same-pulse detections by Alice and Bob are considered for the secure key, thus eliminating all multiphoton events at a cost of lower signal rates.
@@
Figure 4: Experimental results for emulating the DTB protocol. (a) Controlling the photon emission counts by the excitation laser intensity (405 nm, 2 MHz rep. rate, 2 ns pulse duration). Red regions mark $S_{0}$, green regions mark $S_{1}$ and blue regions mark $S_{2}$ appearing in Fig. 1(c). (b) $g^{(2)}(0)$ for the two different intensities, $S_{1}$, $S_{2}$. (c) Calculation of $\gamma$ values as a function of $\eta_{C}$. The blue box represents the $\eta_{C}$ range demonstrated in our devices [34, 36], all with $\gamma>2$ dB.
@@
Figure 5: Experimental results for emulating the HPP protocol. Second-order correlation measurements of the bare gCQD SPS without (a) and with (b) HPP post-processing scheme, resulting in near-zero $g^{(2)}(0)$. The error is mainly due to the SPD dark noise. (c) Calculation showing the optimal transmission $T$, of Alice’s BS, required for maximizing $\gamma$ as a function of $P_{2}$ of the SPS, for two values of $P_{DC}$, corresponding to 100 dark counts per second of the SPD, and SPS excitation repetition rates of $2$ (blue) and $500$ (red) MHz. (d) Calculation of $\gamma$ as a function of $\eta_{C}$ with a fixed $\eta_{D}$ (blue line) and as a function of $\eta_{D}$ with a fixed $\eta_{C}$ (red line). The blue box represents the range demonstrated in our devices [34, 36].
@@
Given a SPS having $\{P_{0},\,P_{1},\,P_{2}\}$, the effective distribution sent to Bob with HPP depends on the reflectance and transmission ($R$ and $T$) of the BS and the detection efficiency of Alice’s detector ($\eta_{D}$), as well as the probability of a dark count at Alice’s detector ($P_{DC}$):
@@
$\tilde{P}_{0}$ $=1-\tilde{P}_{1}-\tilde{P}_{2}$ (11)
Given this new distribution, we can again follow the well established method to estimate the SKR of a BB84 protocol (but now implemented with HPP) [11, 45]. The black line in Fig. 3(a) shows a realistic calculation of the SKR using actual parameters of SPS2 measured in Ref. [34]. A BS reflectance of $50\%$ was chosen. We also used a SPD with $\eta_{D}=0.9$ and $P_{DC}=2\cdot{10}^{-7}$ corresponding to 100 dark counts per second for a $500$ MHz signal rate which are both commonly attainable with current technology. Again, as is seen in the figure, the new BB84+HPP allows for a higher MCL ($\sim 1$ dB) compared to WCS with infinite decoy, due to the extremely low $\tilde{P}_{2}$. Remarkably, again this SKR enhancement can be achieved with SPS far from being ideal having a very low single photon purity.
@@ III.3 Performance Analysis
Fig. 3(b) shows a colormap of the calculated $\gamma$ for different values of $P_{1},P_{2}$ of the SPS (see Supp. Sec. IV). For high $P_{1}$ and sufficiently low $P_{2}$, namely for $P_{1}>1.125P_{2}+0.1927$, there is a region (marked as ”SPS+DTB”) where $\gamma>0\,\text{dB}$, indicating that the use of an SPS with BB84+DTB protocol is advantageous over BB84 including decoy states with WCS (WCS+D). In this region, we highlight in orange dots two known non-classical sources [42, 43] along with our device (SPS1) consisting of a gCQD coupled to a metal-dielectric Bragg nanoantenna [36] (purple), demonstrating that already existing sources, when combined with the DTB protocol, can outperform even ideal WCS protocols. Notably, the use of DTB allows for SPS with a higher probability of two-photon events and shows that bright devices (with fewer vacuum events) can be used for QKD even with single photon purities as low as $\sim 65\%$ and $g^{(2)}(0)$ values as high as $\sim 0.6$.
On the other hand, for SPS with high enough values of $P_{2}$, there is a region where BB84+HPP is advantageous over WCS+D, as shown in the right corner of Fig. 3(b) (marked as ”SPS+HPP”), where $\gamma>0\,\text{dB}$. In the implementation of the HPP protocol, most one-photon events are discarded, and the key is composed largely of two-photon emission events, as indicated in Eq. 9, where the first term is dominant since $P_{DC}\ll P_{1},\,P_{2}$. Therefore, the HPP method is most advantageous in the regime where $P_{2}$ is large, regardless of $P_{1}$. The negligible probability of two-photon events after purification ($<{10}^{-7}$) and minor contributions of the second and third terms in Eq. 9 result in an effectively pure source, differing from a perfect SPS only in brightness (through the zero-photon probability).
As indicated by the first term of Eq. 9, the probability of sending one photon is linearly dependent on $\eta_{D}$, suggesting that the minimum value of $P_{2}$ required for $\gamma>0\,\text{dB}$ is inversely proportional to Alice’s detection efficiency. Calculations yield this relation, giving the condition $P_{2}\geq 0.37/\eta_{D}$ for a balanced 50:50 BS. The values of $\gamma$ and the separation line between the WCS and HPP in Fig. 3(b) are evaluated for the realistic case $\eta_{D}=0.9$, yielding an expected enhancement over WCS when $P_{2}>0.41$.
The black dot in Fig. 3(b) represents an actual SPS device consisting of a gCQD coupled to a Bragg antenna with a plasmonic nanocone (SPS2), demonstrated in Ref. [34], again showing that existing imperfect SPS+HPP can compete with WCS+D.
@@ IV Experimental Results
After showing that existing SPSs, combined with the two new protocols can outperform WCS in terms of MCL, here we demonstrate experimental proof-of-concept emulations of both protocols using our bare single gCQD presented in Fig. 1(c) as our imperfect SPS emulator.
@@
To experimentally demonstrate the feasibility of the DTB scheme, we use three different pulsed excitation intensities generated by a 405 nm diode laser, marked as $S_{0}=0$, $S_{1}$, and $S_{2}$ in Fig. 1(c) (see Supp. Sec. V). Fig. 4(a) demonstrates photon emission control showing stable photon counts under each excitation intensity. This allows an easy control of the photon statistics, as shown in Fig. 4(b), which presents the $g^{(2)}(0)$ values for $S_{1}$ and $S_{2}$, consistent with the modification of photon statistics shown in Fig. 1, where $P^{(S_{1})}_{1}=0.05$, $P^{(S_{2})}_{1}=0.4$ and $P^{(S_{1})}_{2}=0.0005$, $P^{(S_{2})}_{2}=0.15$.
@@
As shown previously, the gCQD SPS with a near unity collection efficiency can outperform WCS+D, if implemented with the DTB protocol. In Fig. 4(c), we numerically calculate the expected $\gamma$ of a gCQD SPS device in terms of $\eta_{C}$ (see Supp. Sec. VI). Remarkably, with the current measured parameters, $\gamma>0\,\text{dB}$ already for $\eta_{C}>0.3$, which is well within reach. As an example, our previously demonstrated gCQD based SPS devices [36] has $\eta_{C}>0.7$ (marked by a blue rectangle) leading to an expected $\gamma>2\,\text{dB}$, a significant improvement over WCS+D.
@@
Moving to emulation of the HPP protocol, we show experimental results of the purification of the gCQD emission in Fig. 5 (see Supp. Sec. V), using a 50:50 BS. The second-order correlation measurements of the gCQD without (Fig. 5(a)) and with (Fig. 5(b)) the HPP post-processing protocol are presented. As can be seen, a near-zero $g^{(2)}(0)$, limited only by detector noise is achieved, competing with state-of-the-art demonstrations [32, 46]. We note that with HPP, the photon rate decreases to $\sim 0.5P_{2}\eta_{D}$, but the collection efficiency $\eta_{C}$ is not affected.
Using the results in Fig. 5(a),(b), we find that $P_{3}\sim 10^{-7}$, therefore we conclude that $P_{N>2}$ is negligible comparing to $P_{1}$ and $P_{2}$ (see Supp. Sec. III), justifying our initial assumptions.
@@
Interestingly, the HPP efficiency can be optimized by adjusting $T$, $R$ of Alice’s BS, depending on $P_{2}$ of the source. In Fig. 5(c) we show a calculation of the optimal beam splitter (BS) transmission $T$ to reach the highest MCL, as a function of $P_{2}$, for two dark-counts probabilities $P_{DC}$. For low $P_{DC}$ the optimal transmittance is roughly $0.5$, however for higher values of $P_{DC}$, the optimal $T$ is smaller than 0.5 and decreases with increasing two-photon probability, in order to minimize the probability for a dark count at Alices’ detector simultaneously with two-photons transmission to Bob.
Lastly, in Fig. 5(d) we plot he calculated $\gamma$ for a gCQD based SPS with the above parameters, as a function of $\eta_{D}$, for a fixed $\eta_{C}=0.9$ (red line), and as a function of $\eta_{C}$ for a fixed $\eta_{D}=0.9$. Again, the blue box, representing demonstrated values of SPS devices based on gCQD coupled to nano-antennas, shows an improvement over WCS+D.

### v2

@@ III.2 Heralded Purification - HP
In Fig. 2(b), we present the experimental configuration of the heralded purification protocol (BB84+HP). In this scheme, we consider an SPS consisting of a gCQD on a hybrid nanocone-antenna device (titled here SPS2). Such a device showed both $\eta_{C}\simeq 1$ together with high attainable values of $P_{2}$, resulting from a large Purcell factor induced by the nanocone which significantly enhances $QY_{XX}$ [41, 52]. Alice operates at a single excitation power $S_{3}>1$, deep in the saturation regime, to excite the BX state with a very high probability, thus maximizing $P_{2}$, which is now only limited by $QY_{XX}$ and $QY_{X}$ [41]. A beam-splitter (BS) and a single photon detector (SPD) are added as a purification stage in Alice’s system, in the emission line of the SPS and before the standard BB84 encoding unit. As $P_{N>2}$ are negligible (see Appendix D), a real photon detection event in Alice’s detector sets $\tilde{P}_{2}=0$ for that pulse, where $\tilde{P}_{n}$ is the effective photon number probability after the purification stage. In the sifting step of the QKD protocol, only events with same-pulse detections by Alice and Bob are considered for the secure key, thus eliminating all multiphoton events at a cost of lower signal rates.
@@
Figure 4: Experimental results for emulating the DTB protocol. (a) Controlling the photon emission counts by the excitation laser intensity (405 nm, 2 MHz rep. rate, 2 ns pulse duration). Red regions mark $S_{0}$, green regions mark $S_{1}$ and purple regions mark $S_{2}$ appearing in Fig. 1(c). The time steps (100 sec) taken for the power modulation of the excitation laser were chosen to emphasize the feasible control of the photon emission. (b) $g^{(2)}(0)$ for the two different intensities, $S_{1}$, $S_{2}$. (c) Calculation of $\gamma$ values as a function of $\eta_{C}$. The blue box represents the $\eta_{C}$ range demonstrated in our devices [41, 43], all with $\gamma>2$ dB.
@@
Figure 5: Experimental results for emulating the HP protocol. Second-order correlation measurements of the bare gCQD SPS without (a) and with (b) the HP post-processing scheme, resulting in near-zero $g^{(2)}(0)$. The error is mainly due to the SPD dark noise. (c) Calculation showing the optimal transmission $T$, of Alice’s BS, required for maximizing $\gamma$ as a function of $P_{2}$ of the SPS, for two values of $P_{DC}$, corresponding to 100 dark counts per second of the SPD, and SPS excitation repetition rates of $2$ (blue) and $500$ (red) MHz. (d) Calculation of $\gamma$ as a function of $\eta_{C}$ with a fixed $\eta_{D}$ (blue line) and as a function of $\eta_{D}$ with a fixed $\eta_{C}$ (red line). The blue box represents the range demonstrated in our devices [41, 43].
@@
Given an SPS having $\{P_{0},\,P_{1},\,P_{2}\}$, the effective distribution sent to Bob with the HP protocol depends on the reflectance and transmission ($R$ and $T$) of the BS and the detection efficiency of Alice’s detector ($\eta_{D}$), as well as the probability of a dark count at Alice’s detector ($P_{DC}$):
@@
Given this new distribution, we can again follow the well established method to estimate the SKR of a BB84 protocol (but now implemented with HP) [11, 53]. Here, the estimated SKR after privacy amplification and error correction, defined as $R_{2}$ for the HP protocol, is given by [2]:
$R_{2}\geq q\cdot Q_{s}\{-H_{2}(E_{s})+\Omega\left[1-H_{2}(\frac{E_{s}}{\Omega})\right]\}$ (10)
where $\Omega=\frac{\tilde{P}_{1}\cdot Y_{1}}{Q_{s}}$ is the relative error of the quantum channel, and $H_{2}$, $q$, $E_{s}$, $Q_{s}$ are defined similarly to the DTB protocol.
The black line in Fig. 3(a) shows a realistic calculation of the SKR using actual parameters of SPS2 measured in Ref. [41] (presented in Table 1 in Appendix G). A BS reflectance of $50\%$ was chosen, and we specifically use $P_{DC}=2\cdot{10}^{-7}$ corresponding to 100 dark counts per second for a $500$ MHz signal rate which are both commonly attainable with current technology. Again, as is seen in the figure, the new BB84+HP allows for a higher MCL ($\sim 1$ dB) compared to WCS with infinite decoy, due to the extremely low $\tilde{P}_{2}$. Remarkably, this SKR enhancement can be achieved with SPS far from being ideal having a very low single photon purity.
@@ III.3 Performance Analysis
Fig. 3(b) shows a colormap of the calculated $\gamma$ for different values of $P_{1},P_{2}$ of the SPS. For high $P_{1}$ and sufficiently low $P_{2}$, namely for $P_{1}>1.125P_{2}+0.1927$, there is a region (marked as ”DTB superiority region”) where $\gamma>0\,\text{dB}$, indicating that the use of an SPS with BB84+DTB protocol is advantageous over WCS with BB84 including infinite decoy states (marked as ”WCS+Decoy”). In this region, we highlight in orange dots two known non-classical sources [49, 50] along with our device (SPS1) consisting of a gCQD coupled to a metal-dielectric Bragg nanoantenna [43] (purple), demonstrating that already existing sources, when combined with the DTB protocol, can outperform even ideal WCS protocols. Notably, the use of DTB allows for an SPS with higher probabilities of two-photon events and shows that bright devices (with fewer vacuum events) can be used for QKD even with single photon purities as low as $\sim 65\%$ and $g^{(2)}(0)$ values as high as $\sim 0.6$.
On the other hand, for an SPS with high enough values of $P_{2}$, there is a region where BB84+HP is advantageous over WCS+Decoy, as shown in the right corner of Fig. 3(b) (marked as ”HP superiority region”), where again $\gamma>0\,\text{dB}$. In the implementation of the HP protocol, most one-photon events are discarded, and the key is composed largely of two-photon emission events, as indicated in Eq. 8, where the first term is dominant since $P_{DC}\ll P_{1},\,P_{2}$. Therefore, the HP method is most advantageous in the regime where $P_{2}$ is large, regardless of $P_{1}$. The negligible probability of two-photon events after purification ($<{10}^{-7}$) and minor contributions of the second and third terms in Eq. 8 result in an effectively pure source, differing from a perfect SPS only in brightness (through the zero-photon probability).
As indicated by the first term of Eq. 8, the probability of sending one photon is linearly dependent on $\eta_{D}$, suggesting that the minimum value of $P_{2}$ required for $\gamma>0\,\text{dB}$ is inversely proportional to Alice’s detection efficiency. Calculations yield this relation, giving the condition $P_{2}\geq 0.37/\eta_{D}$ for a balanced 50:50 BS. The values of $\gamma$ and the separation line between the WCS and HP in Fig. 3(b) are evaluated for the realistic case $\eta_{D}=0.9$, yielding an expected enhancement over WCS when $P_{2}>0.41$. The black dot in Fig. 3(b) represents an actual SPS device consisting of a gCQD coupled to a Bragg antenna with a plasmonic nanocone (SPS2), demonstrated in Ref. [41], again showing that existing imperfect SPS+HP can compete with WCS+Decoy.
Finally, we present experimental results of a bare gCQD sample (shown here in purple and black circles with corresponding error bars), where the photon probabilities were experimentally obtained as explained in Appendix C, and the SKR was extracted separately for each protocol with Eqs. 7,10 respectively. The gCQD, excited at several intensities, exhibits different photon emission statistics for each intensity, thus allowing to demonstrate both the DTB and HP protocols. As seen, even bare gCQDs emitting at room temperature, without any special antennas, can reach a the superior regime over WCS+Decoy.
@@ IV.1 Proof-of-Concept Measurements of Protocols
After theoretically showing that existing SPSs, combined with the two new protocols, can outperform WCS in terms of MCL, here we demonstrate experimental proof-of-concept emulations of both protocols using our bare single gCQD presented in Fig. 1(c) as our imperfect SPS emulator.
@@
To experimentally demonstrate the feasibility of the DTB scheme, we use three different pulsed excitation intensities generated by a 405 nm diode laser, marked as $S_{0}=0$, $S_{1}$, and $S_{2}$ in Fig. 1(c) (see Appendix F for the full experimental information). Fig. 4(a) demonstrates photon emission control showing stable photon counts under each excitation intensity. This allows an easy control of the photon statistics, as shown in Fig. 4(b), which presents the $g^{(2)}(0)$ values for $S_{1}$ and $S_{2}$, consistent with the modification of photon statistics shown in Fig. 1, where $P^{(S_{1})}_{1}=0.05$, $P^{(S_{2})}_{1}=0.4$ and $P^{(S_{1})}_{2}=0.0005$, $P^{(S_{2})}_{2}=0.15$.
@@
As shown previously, the gCQD SPS with a near unity collection efficiency can outperform WCS with decoy states, if implemented with the DTB protocol. In Fig. 4(c), we numerically calculate the expected $\gamma$ of a gCQD SPS device in terms of $\eta_{C}$ (see Appendix I). Remarkably, with the current measured parameters, $\gamma>0\,\text{dB}$ already for $\eta_{C}>0.3$, which is is easily attainable even for bare gCQDs, as we show later. Our previously demonstrated gCQD based SPS devices [43] has $\eta_{C}>0.7$ (marked by a blue rectangle) leading to an expected $\gamma>2\,\text{dB}$, already constituting a significant improvement over WCS+Decoy.
@@
Moving to emulation of the HP protocol, we show experimental results of the purification of the gCQD emission in Fig. 5 using a 50:50 BS. The second-order correlation measurements of the gCQD without (Fig. 5(a)) and with (Fig. 5(b)) the HP post-processing protocol are presented. As can be seen, a near-zero $g^{(2)}(0)$, limited only by detector noise is achieved, competing with state-of-the-art demonstrations [39, 28]. We note that with HP, the photon rate decreases to $\sim 0.5P_{2}\eta_{D}$, but the collection efficiency $\eta_{C}$ is not affected.
Using the results in Fig. 5(a)-(b), one can extract the third-order correlation measurements ($g^{(3)}(0,0)$) [54] to determine $P_{3}$, the probability for a three-photon emission (see Appendix D). Using this method, we find that $P_{3}\sim 10^{-5}$, therefore we conclude that $P_{N>2}$ is negligible compared to $P_{1}$ and $P_{2}$, justifying our initial assumptions.
@@
Interestingly, the HP efficiency can be optimized by adjusting $T$, $R$ of Alice’s BS, depending on the $P_{2}$ of the source. In Fig. 5(c) we show a calculation of the optimal BS transmission $T$ to reach the highest MCL, as a function of $P_{2}$, for two dark-counts probabilities $P_{DC}$. For low $P_{DC}$ the optimal transmittance is roughly $0.5$. However, for higher values of $P_{DC}$, the optimal $T$ is smaller than 0.5 and decreases with increasing two-photon probability, in order to minimize the probability for a dark count at Alices’ detector simultaneously with two-photons transmission to Bob. Lastly, in Fig. 5(d) we plot the calculated $\gamma$ for a gCQD-based SPS with the above parameters, as a function of $\eta_{D}$, for a fixed $\eta_{C}=0.9$, and as a function of $\eta_{C}$ for a fixed $\eta_{D}=0.9$. Again, the blue box, representing demonstrated values of SPS devices based on gCQD coupled to nanoantennas, shows an improvement over WCS+Decoy.
IV.2 Experimental Demonstration of the protocols in BB84-QKD using a gCQD
Figure 6: Polarization-based BB84 QKD measurements of a bare gCQD sample generated with the two protocols. (a) The experimental QKD system sketch: Alice’s setup includes a purification unit with a beam-splitter (BS) and a single photon detector (SPD), and an encryption unit with a linear polarizer (LP) and half-wave plate ($\lambda/2$). The free-space quantum channel includes ND filters to vary the channel loss. Bob’s setup includes the BB84 decryption unit with a $\lambda/2$, a polarizing beam-splitter (PBS) and two SPDs. Tomography mapping results for different excitation laser intensities: $S_{1}$ (b), $S_{2}$ (c), and $S_{3}$ (d), respectively, corresponding to the decoy (green frame) and signal (purple frame) states in the DTB protocol, and $S>1$ intensity (blue frame) state in the HP protocol. (e) The SKR results extracted from the measurements from a bare gCQD sample compared to WCS with infinite decoy. The different measured points along represent the extracted SKR results for several ND filter settings (see full results in Appendix J) using the DTB protocol (purple) and the HP protocol (black). The dashed lines represent the corresponding calculated SKR. The DTB results demonstrate a clear improvement in the maximal channel loss, $\gamma$, over the WCS (red line), while the HP protocol results (dashed black) falls below this due to the small $P_{2}$ values of bare gCQDs without a nanoantenna.
Now, we consider a BB84 demonstration using a bare gCQD as the imperfect SPS and utilizing either DTB or HP protocols.
In Fig. 6(a), we show the QKD system that was used to demonstrate a polarization-based BB84 protocol. The photons emitted from the gCQD sample are collected to Alice’s encryption unit, which randomly define both the qubit value and basis [1, 55]. The photons are then propagated through the quantum channel in free-space, where we introduced ND filters in the optical path to simulate channel loss, before arriving at Bob’s decryption unit. Here, we define the two mutually unbiased bases as ’$+$’ (|$\ket{H},\ket{V}$) and ’$\times$’ (|$\ket{D}$,|$\ket{A}$) [55].
By manipulating the different basis settings of both Alice and Bob, the full tomography mapping can be measured. In Fig. 6(b)-(d), we show the tomography results for three different excitation intensities: $S_{1}$, $S_{2}$, and $S_{3}$ without an ND filter in the quantum channel (for the full set of results, see Appendix J).
From these results, the gain $Q_{i}$ and error rates $E_{i}$ for intensity $i$ are extracted using the following relations:
$Q_{i}=\frac{\text{Total \# of detected qubits}}{\text{Total \# of sent qubits}}$ (11)
$E_{i}=\frac{\text{\# of error qubits}}{\text{Total \# of detected qubits}}$ (12)
The total number of detected qubits is the measured counts at both of Bob’s detectors at a given basis, and the total number of error qubits is the number of counts measured on the wrong detector (assuming Alice and Bob randomly chose the same basis, according to the BB84 protocol). The total number of qubits sent by Alice is estimated by the following:
Total # of sent $\text{Total \# of sent qubits}=N\cdot\eta_{A}\cdot\eta_{C_{NA}}$ (13)
where $N$ is the repetition rate of the excitation laser (2 MHz, see Appendix F), $\eta_{A}\simeq 19.5\%$ is the transmission of Alice’s setup measured in our experiments, and $\eta_{C_{NA}}\simeq 14.18\%$ is the transmission of the photon collection optics from the gCQD to Alice extracted previously in Ref. [40].
The $Q_{i}$, $E_{i}$, along with the measured photon statistics $\{P_{0},P_{1},P_{2}\}$ for each intensity $i$, were inserted into Eq. III.1 to extract $Y_{n}$ and $e_{n}$, which were then used for calculating the SKR in Eqs. 7,10 for the DTB and HP protocols respectively. The DTB model requires two decoys ($S_{0}$ and $S_{1}$), where $S_{0}$ is the vacuum state with known parameters (see Appendix G). Here, $S_{2}$ is used as the signal state for the DTB protocol, and $S_{3}$ is used for the HP protocol.
In Fig. 6(c), we plot the SKR of the gCQD measured with the DTB protocol and the HP protocol, for several channel losses set with various ND filters, compared to WCS with infinite decoy states. A clear improvement of $\sim 2$ dB is demonstrated with the DTB protocol, which agree well with the theoretical curve. As expected, the results with the HP protocol do not show an improvement compared to WCS with decoy, due to the low $P_{2}$ in bare gCQD as compared to those obtained in a full SPS2 devices [41], nonetheless, showing the feasibility of the two enhanced-QKD protocols with already available SPSs, and their advantage over the best existing WCS solutions.